# Globalization of perturbative Chern-Simons theory on the moduli space of flat connections in the BV formalism

### v1

@@ 5.3. Extended smoothness assumption
(209) $\widetilde{Z}=\check{Z}e^{\frac{i}{\hbar}\langle[\delta A],\mathsf{a}\rangle}\Big|_{A^{\prime}=A,\delta A^{\prime}=\delta A\mathrm{\,harmonic},\;\delta g=0}=\widehat{Z}\Big|_{A^{\prime}=A,\delta A^{\prime}=\delta A\mathrm{\,harmonic}}$
with $\check{Z}$ as in (171) and $\widehat{Z}$ as in (138), with the r.h.s. considered modulo gauge transformations. Then, assuming 1-extended smoothness, the dQME (175) yields (208).
@@
Step 2. (Building a chain contraction for $\nabla^{G}-i\hbar\Delta_{\mathsf{a}}$.) The cohomology of the complex $\Omega^{\bullet}(\mathcal{M}^{\prime},\mathrm{Dens}^{\frac{1}{2},\mathrm{formal}}(H^{\bullet}_{A}[1])),\nabla^{G}$ is concentrated in form degree 0 and is isomorphic to global half-densities $\mathrm{Dens}^{\frac{1}{2}}(T^{*}[-1]\mathcal{M}^{\prime})$. [^53] See [BCM12, Section 2]. More precisely, one has SDR data $(\mathsf{i},\mathsf{p},\mathsf{K})$ with inclusion $\mathsf{i}=T\underline{\varphi}^{*}$ and projection $\mathsf{p}$ given by evaluating a formal half-density at $\mathsf{a}^{1}=0$,
@@
Next, deform the differential on $\Omega^{\bullet}(\mathcal{M}^{\prime},\mathrm{Dens}^{\frac{1}{2},\mathrm{formal}}(H^{\bullet}_{A}[1]))$ from $\nabla^{G}$ to $\nabla^{G}-i\hbar\Delta_{\mathsf{a}}$. By homological perturbation lemma (Lemma A.2), one has deformed SDR data
(211) $(\mathsf{i}^{\prime},\mathsf{p}^{\prime},\mathsf{K}^{\prime})\colon(\Omega^{\bullet}(\mathcal{M}^{\prime},\mathrm{Dens}^{\frac{1}{2},\mathrm{formal}}(H^{\bullet}_{A}[1])),\nabla^{G}+i\hbar\Delta_{\mathsf{a}})\leadsto(\mathrm{Dens}^{\frac{1}{2}}(T^{*}[-1]\mathcal{M}^{\prime}),\delta).$
@@
Step 3. Using the fact that $K^{\prime}$ defined above satisfies the chain homotopy property $\mathrm{id}=\mathsf{i}\mathsf{p}^{\prime}+[\nabla^{G}-i\hbar\Delta_{\mathsf{a}},\mathsf{K}^{\prime}]$, we have
@@
Denote the first term on the r.h.s. by $Z^{\mathrm{mod}}\colon=\mathsf{i}\mathsf{p}^{\prime}\widetilde{Z}$. Since it is in the image of $\mathsf{i}$, it is $\nabla^{G}$-closed, and hence a global object. Restricting (212) to form degree zero along $\mathcal{M}^{\prime}$ and denoting
(213) $\rho=(\mathsf{K}^{\prime}\widetilde{Z})|_{\Omega^{0}(\mathcal{M}^{\prime},\cdots)},$
@@
Remark 5.17.
Note that 1-extended smoothness implies $\rho=\underline{Z}\cdot O(\hbar^{0})$ and $Z^{\mathrm{mod}}=\underline{Z}\cdot(1+O(\hbar))$.
@@
In the notations of the proof of Theorem 5.16, we have $Z^{\mathrm{glob}}=\mathsf{p}^{\prime}\widetilde{Z}$. Remark 5.17 above implies
@@
(220) $Z^{\mathrm{glob}}(A)=\int_{H^{2}_{A}[-1]\oplus H^{1}_{A}[1]}\mathcal{D}\mathsf{a}^{2}\,\mathcal{D}\zeta\int_{\mathcal{L}=\Omega_{d_{A}^{*}\mathrm{-ex}}[1]}\mathcal{D}\alpha_{\mathrm{fl}}\\ \exp\frac{i}{\hbar}\Bigg(S_{CS}(A+i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}})+\langle\mathsf{a}^{2},\zeta\rangle\\ +\int_{M}\frac{1}{2}\langle\alpha_{\mathrm{fl}},d_{A}G\mathrm{ad}^{*}_{i(\zeta)}\alpha_{\mathrm{fl}}\rangle+\langle\alpha_{\mathrm{fl}},d_{A}G\mathrm{ad}^{*}_{i(\zeta)}i(\mathsf{a}^{2})\rangle\Bigg).$
@@
Figure 12. Example of a Feynman graph for $Z^{\mathrm{glob}}$ (220). Dashed edges correspond to $\mathsf{a}^{2}$$\zeta$ propagators; circle vertices correspond to (221). Selection rules: $\leq 2$ circle vertices on a solid edge, $\leq\dim\mathcal{M}^{\prime}$ dashed edges in total. A solid edge not incident to Chern-Simons cubic vertices should have exactly two circle vertices (as in the top part of the picture).
@@ 5.5. Metric dependence of the global partition function
In this section we are assuming metric-extended smoothness, cf. Definition 5.15.
@@
Consider the following object:
(224) $\overline{Z}^{\mathrm{ren}}\colon=e^{\frac{i}{\hbar}c(\hbar)\frac{S_{\mathrm{grav}}(g,\phi)}{2\pi}}\cdot\pi_{*}\Big(\check{Z}\cdot e^{-\frac{i}{\hbar}\langle[\delta A],\mathsf{a}\rangle}\Big|_{A=A^{\prime},\delta A=\delta A^{\prime}\mathrm{\;harmonic}}\Big)\\ \in\Omega^{\bullet}(\mathrm{Met}\times\mathcal{M}^{\prime},\mathrm{Dens}^{\frac{1}{2},\mathrm{formal}}(H_{A}[1]))$
where $\check{Z}$ is as in (170) and $\pi_{*}$ stands for the passage to the quotient by gauge transformations, cf. 201.
Let
@@
First note that, by restricting (175) to the diagonal $A=A^{\prime}$ and setting $\delta A=\delta A^{\prime}$ to be harmonic, one obtains, under the metric-extended smoothness assumption, the equation
(229) $(\delta_{g}+\nabla^{G}-i\hbar\Delta_{\mathsf{a}})\overline{Z}^{\mathrm{ren}}=0.$
Next, consider the contraction (211). Note that, by metric-extended smoothness assumption, all maps involved do not depend on the metric. Tensoring (211) with the de Rham complex of $\mathrm{Met}$, we obtain the contraction
(230) $(\mathsf{i}^{\prime},\mathsf{p}^{\prime},\mathsf{K}^{\prime})\colon(\Omega^{\bullet}(\mathrm{Met}\times\mathcal{M}^{\prime},\mathrm{Dens}^{\frac{1}{2},\mathrm{formal}}(H_{A}[1])),\delta_{g}+\nabla^{G}-i\hbar\Delta_{\mathsf{a}})\\ \leadsto(\Omega^{\bullet}(\mathrm{Met})\otimes\mathrm{Dens}^{\frac{1}{2}}(T^{*}[-1]\mathcal{M}^{\prime}),\delta_{g}-i\hbar\Delta).$
Since $\mathsf{p}^{\prime}$ is a chain map, it sends the cocycle $\overline{Z}^{\mathrm{ren}}$ of the complex upstairs to a cocycle $\check{Z}^{\mathrm{glob,ren}}$ of the complex downstairs. This proves (226).
@@
(231) $\check{Z}^{\mathrm{glob,ren}}(A,\mathsf{a}^{2})=e^{\frac{i}{\hbar}c(\hbar)\frac{S_{\mathrm{grav}}(g,\phi)}{2\pi}}\cdot\int_{H^{2}_{A}[-1]\oplus H^{1}_{A}[1]}\mathcal{D}\mathsf{a}^{2}\,\mathcal{D}\zeta\int_{\mathcal{L}=\Omega_{d_{A}^{*}\mathrm{-ex}}[1]}\mathcal{D}\alpha_{\mathrm{fl}}\\ \exp\frac{i}{\hbar}\Big(S_{CS}(A+i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}})+\langle\mathsf{a}^{2},\zeta\rangle\\ +\int_{M}\underbrace{\frac{1}{2}\langle\alpha_{\mathrm{fl}},d_{A}G\mathrm{ad}^{*}_{i(\zeta)}\alpha_{\mathrm{fl}}\rangle+\langle\alpha_{\mathrm{fl}},d_{A}G\mathrm{ad}^{*}_{i(\zeta)}i(\mathsf{a}^{2})\rangle}_{-\frac{1}{2}\langle i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}},H_{\delta A^{\prime}=i(\zeta)}(i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}})\rangle}\underbrace{-\frac{1}{2}\langle\alpha_{\mathrm{fl}},\lambda_{\delta g}\alpha_{\mathrm{fl}}\rangle-\langle\alpha_{\mathrm{fl}},\lambda_{\delta g}i(\mathsf{a}^{2})\rangle}_{-\frac{1}{2}\langle i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}},H_{\delta g}(i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}})\rangle}\Big)$
@@
(232) Zˇglob,ren(1)=eic()Sgrav(g,ϕ)2πeiSCS(A)eπi4ψA,gτA1/2([Uncaptioned image]+[Uncaptioned image]+
+[Uncaptioned image]+

### v2

@@ 5.2. A formula for the Grothendieck connection
(222) $0=\pi_{*}I^{*}e^{-\frac{i}{\hbar}\Xi}\underbrace{(\nabla^{\mathcal{D}}-i\hbar\Delta_{\mathsf{a}}-\frac{i}{\hbar}\frac{1}{2}\langle\mathsf{a},F_{\nabla^{\mathbb{H}}}\mathsf{a}\rangle)}_{\nabla^{G-M}}(e^{\frac{i}{\hbar}\Xi}(e^{-\frac{i}{\hbar}\Xi}\check{Z}^{\mathrm{ren}}))\\ =\Big(\nabla^{\mathbb{H}}_{\mathrm{Met}\times\mathcal{M}^{\prime}}+\{\xi^{\mathrm{tot}},-\}+\Delta_{\mathsf{a}}\xi^{\mathrm{tot}}-i\hbar\Delta_{\mathsf{a}}+\\ +\frac{i}{\hbar}(\nabla^{\mathbb{H}}_{\mathrm{Met}\times\mathcal{M}^{\prime}}\xi^{\mathrm{tot}}+\frac{1}{2}\{\xi^{\mathrm{tot}},\xi^{\mathrm{tot}}\}-\frac{1}{2}\langle\mathsf{a},F_{\nabla^{\mathbb{H}}}\mathsf{a}\rangle+\underbrace{\frac{1}{2}\langle[\mathsf{a},\mathsf{a}],KHi[\delta A]\rangle}_{Y})\Big)\overline{Z}^{\mathrm{ren}}.$
Here the differential operator in brackets in lines 2 and 3 – except the term $Y$ – arises as a conjugation of the Gauss-Manin superconnection by $e^{\frac{i}{\hbar}\Xi}$, restricted to $A^{\prime}=A$ and $\delta A^{\prime}=\delta A$ harmonic and reduced modulo gauge transformations.
Next we comment on the $Y$ term appearing in (222). First consider for simplicity the case of $g$ fixed. In fact, $I^{*}(\nabla^{G-M}\check{Z}^{\mathrm{ren}})$ cannot be computed a priori as a differential operator (denote it $I^{*}\nabla^{G-M}$) acting on $I^{*}\check{Z}^{\mathrm{ren}}$. The reason is that when we move infinitesimally on $\mathrm{FC}^{\prime}\stackrel{{\scriptstyle\mathrm{Diag}}}{{\hookrightarrow}}\mathrm{FC}^{\prime}\times\mathrm{FC}^{\prime}$, the notion of harmonic forms changes, so, in order to evaluate $I^{*}(\nabla^{G-M}\check{Z}^{\mathrm{ren}})$ on a tuple of harmonic tangent vectors, it is insufficient to know $I^{*}(\check{Z}^{\mathrm{ren}})$ – one needs also the value of $\check{Z}^{\mathrm{ren}}$ on a collection of harmonic tangent vectors and one non-harmonic (arising as a variation of a harmonic tangent vector). More precisely: given harmonic tangent vectors $\alpha_{0},\ldots,\alpha_{p}\in T_{(A,A)}\mathrm{FC}^{\prime}_{\mathrm{Diag}}$, we extend them to harmonic vector fields $\widetilde{\alpha}_{0},\ldots,\widetilde{\alpha}_{p}$ on $\mathrm{FC}^{\prime}_{\mathrm{Diag}}$. Then we have
(223) $(\nabla^{\mathcal{D}}\check{Z}^{\mathrm{ren}})(\widetilde{\alpha}_{0},\ldots,\widetilde{\alpha}_{p})=\sum_{i=0}^{p}\nabla^{\mathcal{D}}_{\widetilde{\alpha}_{i}}\check{Z}^{\mathrm{ren}}(\widetilde{\alpha}_{0},\ldots\widehat{\widetilde{\alpha}_{i}}\ldots,\widetilde{\alpha}_{p})\\ -\sum_{i,j=0}^{p}\frac{1}{2}\check{Z}^{\mathrm{ren}}([\widetilde{\alpha}_{i},\widetilde{\alpha}_{j}],\widetilde{\alpha}_{0},\ldots\widehat{\widetilde{\alpha}_{i}}\ldots\widehat{\widetilde{\alpha}_{j}}\ldots,\widetilde{\alpha}_{p}).$
Then, choosing the extensions $\widetilde{\alpha}_{i}|_{A+\epsilon\beta}=\alpha_{i}-\epsilon(K\mathrm{ad}_{\beta}\alpha_{i}+dG\mathrm{ad}^{*}_{\beta}\alpha_{i})+O(\epsilon^{2})$ (cf. (82), (83)), we have $\widetilde{\alpha}_{i},\widetilde{\alpha}_{j}]|_{(A,A)}=2dG*[\alpha_{i},*\alpha_{j}]$. Hence, we have
(224) $I^{*}\nabla^{G-M}\check{Z}^{\mathrm{ren}}=(I^{*}\nabla^{G-M})(I^{*}\check{Z}^{\mathrm{ren}})+I^{*}(\langle-dG\mathrm{ad}^{*}_{\delta A}\delta A,\frac{\delta}{\delta(\delta A)}\rangle\check{Z}^{\mathrm{ren}})$
Here the two terms on the right corresponds to the two terms on the right in (223). Allowing $g$ to vary (and so allowing the tangent vectors to have component along $T_{g}\mathrm{Met}$), the obvious extension of this computation yields
(225) $I^{*}\nabla^{G-M}\check{Z}^{\mathrm{ren}}=(I^{*}\nabla^{G-M})(I^{*}\check{Z}^{\mathrm{ren}})+I^{*}(\langle dK\underbrace{H}_{H_{\delta A^{\prime}}+H_{\delta g}}\delta A,\frac{\delta}{\delta(\delta A)}\rangle\check{Z}^{\mathrm{ren}})$
In terms of the Feynman diagram expansion of $\check{Z}$ of Remark 4.22, the rightmost term in (225) has the following contributions:
  1. (i)
    $Hi[\delta A]$ replacing a gray vertex in a graph (from the derivation acting on $\delta A$ in a gray vertex),
  2. (ii)
    replacing a white vertex in a graph by $\mathrm{ad}_{KHi[\delta A]}$ (from the derivation acting on $\delta A^{\prime}=\delta A$ in a white vertex).
Contributions (i) and (ii) above mutually cancel, except for one special graph in (i) – a cubic corolla with one leaf attached to a gray vertex, replacing the latter with $Hi[\delta A]$ as in (i) above, see Figure 12. This yields $Y\check{Z}^{\mathrm{ren}}$, resulting eventually in the $Y$ term correction in (222).
Figure 12. Cancellation mechanism in the rightmost term in (225). Circle vertex with a cross stands for either white or black vertex. Dashed region represents some graph.
Now that (222) is proven, we show that it can be simplified. Consider the dQME (175) on $\check{Z}^{\mathrm{ren}}$, restrict it to lowest order in $\hbar$ (tree diagram contributions) and to 2-forms on $\underline{\mathcal{U}}$. Then, reducing to $\mathrm{Met}\times\mathcal{M}^{\prime}$ by applying $\pi_{*}I^{*}$, we obtain the equation
(226) $\nabla^{\mathbb{H}}_{\mathrm{Met}\times\mathcal{M}^{\prime}}\xi^{\mathrm{tot}}+\frac{1}{2}\{\xi^{\mathrm{tot}},\xi^{\mathrm{tot}}\}-\frac{1}{2}\langle\mathsf{a},F_{\nabla^{\mathbb{H}}}\mathsf{a}\rangle+Y=0$
where the $Y$ term appears by the same mechanism as above. Hence, (222) simplifies to
(227) $(\delta_{g}+\underbrace{\nabla^{\mathbb{H}}_{\mathcal{M}}+\{\xi,-\}+\Delta_{\mathsf{a}}\xi}_{\nabla^{G}}+\{\mu,-\}+\Delta_{\mathsf{a}}\mu-i\hbar\Delta_{\mathsf{a}})\overline{Z}^{\mathrm{ren}}=0,$
which proves (219).
Finally, (221) is an immediate consequence of (219) by restricting to a $g$-fixed slice of $\mathrm{Met}\times\mathcal{M}^{\prime}$. ∎
Proof of Theorem 5.10.
@@
Step 1. (Building a chain contraction for $\nabla^{G}-i\hbar\Delta_{\mathsf{a}}$.) The cohomology of the complex $(\Omega^{\bullet}(\mathcal{M}^{\prime},\mathcal{D}),\nabla^{G})$ is concentrated in form degree 0 and is isomorphic to global half-densities $\mathrm{Dens}^{\frac{1}{2}}(T^{*}[-1]\mathcal{M}^{\prime})$. [^55] See [BCM12, Section 2]. More precisely, one has SDR data $(\mathsf{i},\mathsf{p},\mathsf{K})$ with inclusion $\mathsf{i}=T\underline{\varphi}^{*}$ and projection $\mathsf{p}$ given by evaluating a formal half-density at $\mathsf{a}^{1}=0$,
@@
Next, deform the differential on $\Omega^{\bullet}(\mathcal{M}^{\prime},\mathcal{D})$ from $\nabla^{G}$ to $\nabla^{G}-i\hbar\Delta_{\mathsf{a}}$. By homological perturbation lemma (Lemma A.2), one has deformed SDR data
(229) $(\mathsf{i}^{\prime},\mathsf{p}^{\prime},\mathsf{K}^{\prime})\colon(\Omega^{\bullet}(\mathcal{M}^{\prime},\mathcal{D}),\nabla^{G}-i\hbar\Delta_{\mathsf{a}})\leadsto(\mathrm{Dens}^{\frac{1}{2}}(T^{*}[-1]\mathcal{M}^{\prime}),\delta).$
@@
Step 2. Using the fact that $K^{\prime}$ defined above satisfies the chain homotopy property $\mathrm{id}=\mathsf{i}\mathsf{p}^{\prime}+[\nabla^{G}-i\hbar\Delta_{\mathsf{a}},\mathsf{K}^{\prime}]$, we have
@@
where $\widetilde{Z}$ is as in (220). Denote the first term on the r.h.s. by $Z^{\mathrm{mod}}\colon=\mathsf{i}\mathsf{p}^{\prime}\widetilde{Z}$. Since it is in the image of $\mathsf{i}$, it is $\nabla^{G}$-closed, and hence a global object. Restricting (230) to form degree zero along $\mathcal{M}^{\prime}$ and denoting
(231) $\rho=(\mathsf{K}^{\prime}\widetilde{Z})|_{\Omega^{0}(\mathcal{M}^{\prime},\mathcal{D})},$
@@
Lemma 5.12.
We have the following ansatz for $\hbar$-dependence of $\rho$ and $Z^{\mathrm{mod}}$:
(232) $\rho$ $=$ $\underline{Z}\cdot f(\hbar),$
(233) $Z^{\mathrm{mod}}$ $=$ $\underline{Z}\cdot(1+\hbar\,g(\hbar)),$
where $f,g\in C^{\infty}(\mathcal{M}^{\prime},\underbrace{\widehat{\mathrm{Sym}}(H_{A}[1])^{*}}_{\mathbb{F}})[[\hbar]]$ are some formal power series in $\hbar$ and $\mathsf{a}$.
Proof.
Both properties follow from the fact that
(234) $\widetilde{Z}=\underbrace{e^{\frac{i}{\hbar}S_{CS}(A)}\tau_{A}^{\frac{1}{2}}e^{\frac{\pi i}{4}\psi_{A}}}_{\underline{Z}_{0}}\,e^{\frac{i}{\hbar}\omega+h}$
with $\omega\in\Omega^{\geq 2}(\mathcal{M}^{\prime},\mathbb{F})$ independent of $\hbar$ and with $h\in\Omega^{\geq 0}(\mathcal{M}^{\prime},\mathbb{F})[[\hbar]]$. Hence, the $p$-form component of $\widetilde{Z}$ along $\mathcal{M}^{\prime}$ is
(235) $\widetilde{Z}^{(p)}=\underline{Z}_{0}\cdot\hbar^{-\left[\frac{p}{2}\right]}F^{(p)}=\underline{Z}\cdot\hbar^{-\left[\frac{p}{2}\right]}F^{{}^{\prime}(p)}$
with $F^{(p)},F^{{}^{\prime}(p)}\in\Omega^{p}(\mathcal{M}^{\prime},\mathbb{F})[[\hbar]]$. Therefore,
(236) $\rho=\mathsf{K}^{\prime}\widetilde{Z}\big|_{\Omega^{0}(\mathcal{M}^{\prime},\mathcal{D})}=\sum_{k\geq 0}(\mathsf{K}i\hbar\Delta_{\mathsf{a}})^{k}\mathsf{K}\left(\widetilde{Z}|_{\Omega^{k+1}(\mathcal{M}^{\prime},\mathcal{D})}\right)=\underline{Z}_{0}\cdot\sum_{k\geq 0}\hbar^{k-\left[\frac{k+1}{2}\right]}G_{k}$
with $G_{k}\in\Omega^{0}(\mathcal{M}^{\prime},\mathbb{F})[[\hbar]]$. This implies (232) and – using (215) – implies also (233). ∎
@@
In the notations of the proof of Theorem 5.10, we have
(238) $Z^{\mathrm{glob}}=\mathsf{p}^{\prime}\widetilde{Z}$.
Lemma 5.12 above implies
@@
(244) $Z^{\mathrm{glob}}(A)=\mathrm{RT}\int_{H^{2}_{A}[-1]\oplus H^{1}_{A}[1]}\mathcal{D}\mathsf{a}^{2}\,\mathcal{D}\zeta\int_{\mathcal{L}=\Omega_{d_{A}^{*}\mathrm{-ex}}[1]}\mathcal{D}\alpha_{\mathrm{fl}}\\ \exp\frac{i}{\hbar}\Bigg(S_{CS}(A+i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}})+\langle\mathsf{a}^{2},\zeta\rangle\\ +\int_{M}\frac{1}{2}\langle\alpha_{\mathrm{fl}},d_{A}G\mathrm{ad}^{*}_{i(\zeta)}\alpha_{\mathrm{fl}}\rangle+\langle\alpha_{\mathrm{fl}},d_{A}G\mathrm{ad}^{*}_{i(\zeta)}i(\mathsf{a}^{2})\rangle\Bigg).$
@@
Figure 13. Example of a Feynman graph for $Z^{\mathrm{glob}}$ (244). Dashed edges correspond to $\mathsf{a}^{2}$$\zeta$ propagators; white vertices correspond to (245). Selection rules: $\leq 2$ white vertices on a solid edge, $\leq\dim\mathcal{M}^{\prime}$ dashed edges in total. A solid edge not incident to Chern-Simons cubic vertices should have exactly two white vertices (as in the top part of the picture).
The symbol $\mathrm{RT}$ “remove trees” in (244) stands for removing from the perturbative expansion Feynman graphs with the property that if one cuts all dashed edges ($\mathsf{a}^{2}-\zeta$ propagators), one of the remaining connected components is a tree with exactly one $H$-vertex.
@@ 5.5. Metric dependence of the global partition function
Consider $\overline{Z}^{\mathrm{ren}}$ defined by (218) and let
@@
First, recall from Lemma 5.11 that $\overline{Z}^{\mathrm{ren}}$ satisfies
(252) $(\delta_{g}+\nabla^{G}+\{\mu,-\}+\Delta_{\mathsf{a}}\mu-i\hbar\Delta_{\mathsf{a}})=0.$
Next, consider the differential perturbation of the contraction
(253) $(\mathsf{i},\mathsf{p},\mathsf{K})\colon(\Omega^{\bullet}(\mathrm{Met}\times\mathcal{M}^{\prime},\mathcal{D},\nabla^{G})\\ \leadsto(\Omega^{\bullet}(\mathrm{Met})\otimes\mathrm{Dens}^{\frac{1}{2}}(T^{*}[-1]\mathcal{M}^{\prime}),\mathrm{zero\;differential}),$
deforming $\nabla^{G}\rightarrow\nabla^{G}+\underbrace{\delta_{g}+\{\mu,-\}+\Delta_{\mathsf{a}}\mu-i\hbar\Delta_{\mathsf{a}}}_{\varkappa}$. By homological perturbation lemma, we obtain the deformed contraction
(254) $(\mathsf{i}^{\prime\prime},\mathsf{p}^{\prime\prime}=\mathsf{p}^{\prime},\mathsf{K}^{\prime\prime})\colon(\Omega^{\bullet}(\mathrm{Met}\times\mathcal{M}^{\prime},\mathcal{D}),\nabla^{G}+\varkappa)\\ \leadsto(\Omega^{\bullet}(\mathrm{Met})\otimes\mathrm{Dens}^{\frac{1}{2}}(T^{*}[-1]\mathcal{M}^{\prime}),\delta_{g}-i\hbar\Delta).$
We remark that:
  1. (i)
    The induced differential in (254) is $\sum_{k\geq 0}\mathsf{p}\varkappa(-\mathsf{K}\varkappa)^{k}\mathsf{i}$. Since the image of $\mathsf{i}$ is in zero-forms on $\mathcal{M}^{\prime}$, $K$ reduces the form degree on $\mathcal{M}^{\prime}$ by one and $\varkappa$ does not decrease the form degree on $\mathcal{M}^{\prime}$, all terms with $k>0$ vanish and the induced differential is $\mathsf{p}\varkappa\mathsf{i}=\delta_{g}-i\hbar\Delta$.
  2. (ii)
    One has $\mathsf{p}^{\prime\prime}=\mathsf{p^{\prime}}$ – the chain projection in (229) extended trivially along $\mathrm{Met}$. Indeed, one has
    (255) $\mathsf{p}^{\prime\prime}=\sum_{k\geq 0}\mathsf{p}(-\varkappa\mathsf{K})^{k}.$
    Since $\mathsf{p}$ evaluates the constant term in $\mathsf{a}^{1}$, $\mathsf{K}$ increases the degree in $\mathsf{a}^{1}$ and $\varkappa=-i\hbar\Delta_{\mathsf{a}}$ plus terms that do not decrease the degree in $\mathsf{a}^{1}$, in (255) one can replace $\varkappa$ with $i\hbar\Delta_{\mathsf{a}}$, leading to $\mathsf{p}^{\prime\prime}=\mathsf{p}^{\prime}$.
Since $\mathsf{p}^{\prime}$ is a chain map, it sends the cocycle $\overline{Z}^{\mathrm{ren}}$ of the complex upstairs in (254) to a cocycle $\check{Z}^{\mathrm{glob,ren}}$ of the complex downstairs. This proves (249).
@@
(256) $\check{Z}^{\mathrm{glob,ren}}(A,\mathsf{a}^{2})=e^{\frac{i}{\hbar}c(\hbar)\frac{S_{\mathrm{grav}}(g,\phi)}{2\pi}}\cdot\mathrm{RT}\int_{H^{2}_{A}[-1]\oplus H^{1}_{A}[1]}\mathcal{D}\mathsf{a}^{2}\,\mathcal{D}\zeta\int_{\mathcal{L}=\Omega_{d_{A}^{*}\mathrm{-ex}}[1]}\mathcal{D}\alpha_{\mathrm{fl}}\\ \exp\frac{i}{\hbar}\Big(S_{CS}(A+i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}})+\langle\mathsf{a}^{2},\zeta\rangle\\ +\int_{M}\underbrace{\frac{1}{2}\langle\alpha_{\mathrm{fl}},d_{A}G\mathrm{ad}^{*}_{i(\zeta)}\alpha_{\mathrm{fl}}\rangle+\langle\alpha_{\mathrm{fl}},d_{A}G\mathrm{ad}^{*}_{i(\zeta)}i(\mathsf{a}^{2})\rangle}_{-\frac{1}{2}\langle i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}},H_{\delta A^{\prime}=i(\zeta)}(i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}})\rangle}\underbrace{-\frac{1}{2}\langle\alpha_{\mathrm{fl}},\lambda_{\delta g}\alpha_{\mathrm{fl}}\rangle-\langle\alpha_{\mathrm{fl}},\lambda_{\delta g}i(\mathsf{a}^{2})\rangle}_{-\frac{1}{2}\langle i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}},H_{\delta g}(i(\mathsf{a}^{2})+\alpha_{\mathrm{fl}})\rangle}\Big)$
@@
The symbol $\mathrm{RT}$ in (256) stands for removing Feynman graphs such that if all dashed edges are deleted, there is a connected component which is a tree with one white vertex or one black vertex.